Cos 240 + tan 120 + sin 510 + tan 210 = 0 prove this 

Grade:11

1 Answers

Harshit Singh
3 years ago
Dear Student

cos 240 = – ½
tan 120 = –√3
sin 510 = 1/2
tan 210 = 1/ √3

Adding all these values
Cos 240 + tan 120 + sin 510 + tan 210= 1/√3 –√3
Please check the question again

Thanks